OpenStudy (anonymous):

Graph the two equations to solve the system. What is the ordered pair solution? x + y = 6 and x - y = 4

hero (hero):

Add both equations together x + y = 6 x - y = 4 _________ 2x = 10 Solve for x, then find the value for y.

OpenStudy (anonymous):

Is y 4?

hero (hero):

Did you find x? You have to find that first.

hero (hero):

By the way, y is not 4. I have no idea how you got that.

OpenStudy (anonymous):

yes ... x is 4 and y is 2 ... Sorry got them mixed up @Hero

hero (hero):

That's not correct.

hero (hero):

Explain to me something.....What method did you use to find x. Show me your steps. I get the feeling that you guessed.

OpenStudy (anonymous):

Wait I see what I did wrong ... x is 6 right ?

hero (hero):

You're guessing again. It almost seems as if there's not hope for you.

OpenStudy (anonymous):

Ugh ,... I don't know ... I am not guessing but, I don't know how to get there ...

hero (hero):

2x = 10 x = 10/2 x = 5 y =1 5 + 1 = 6 5 - 1 = 4